# Finding Safety Neurons in Large Language Models

### v1

@@ E.3 More Alignment Tax Results
Spearman’s rank correlation coefficients between different preference neurons of Mistral-7b and Gemma-7b are shown in Figure 12. For Mistral-7b, we observe results similar to Llama2-7b. However, Gemma-7b shows anomalies when aligned on RewardBench-Safety, which we attribute to the small dataset size (~1k samples) compared to the larger number of neurons Gemma-7b. This discrepancy likely leads to insufficient training. However, this discrepancy does not affect our explanation of the alignment tax (Table 3).
E.4 More Safeguard Results
Method JL HB HH RT Avg
Activations from 150 neurons
Safety Neuron $\mathbf{78.2}$ $\mathbf{76.0}$ $77.8$ $\mathbf{79.1}$ $\mathbf{77.8}$
Random Neuron $74.4$ $69.2$ $\mathbf{78.0}$ $78.6$ $75.0$
Random Neuron $76.4$ $69.0$ $77.8$ $78.4$ $75.4$
Majority $54.6$ $62.0$ $63.6$ $65.8$ $65.8$
Activations from 1500 neurons
Safety Neuron $80.5$ $\mathbf{72.5}$ $82.8$ $81.6$ $\mathbf{79.4}$
Random Neuron $\mathbf{80.8}$ $69.1$ $\mathbf{83.4}$ $\mathbf{82.2}$ $78.8$
Random Neuron $79.9$ $69.8$ $82.6$ $81.7$ $78.5$
Majority $54.6$ $62.0$ $63.6$ $65.8$ $65.8$
Table 8: Accuracy (%) of logistic regression classifier trained on Llama2-7b neuron activations collected on Beavertails dataset. JL = JailBreakLLMs, HB = HarmBench, HH = HH-Harmless, RT = RedTeam. Random Neuron refers to randomly sampled neurons with each layer’s neuron count matching that of safety neurons, Random Neuron denotes neurons randomly sampled from the last layer, and Majority denotes the classifier that always predicts the majority class in labels. All the random neurons are sampled 5 times with different random seeds and report the average results.
@@ E.4 More Safeguard Results
To validate the generalization ability of these neuron activations, we use activations from the Beavertails as the training set and others as the test set, training a simple logistic regression classifier and using accuracy as our metric. In addition to safety neurons, we employ neurons identified through other strategies as baselines, including (1) Random Neuron, which refers to randomly sampled neurons with each layer’s neuron count matching that of safety neurons; (2) Random Neuron (last), which denotes neurons randomly sampled entirely from the last layer, based on the hypothesis that the last layer’s neurons directly affect the model’s output, making this a potentially strong baseline; (3) Majority, which is a classifier that always predicts the majority class in labels to account for the potential impact of class imbalance in the dataset and ensure the model’s true performance is reflected. For all experiments requiring randomly sampled neurons, we repeat the process 5 times using different random seeds and report the averaged results.
@@
We train and test the classifier using activations from different numbers of neurons, as shown in Figure 11. The results indicate that the test accuracy almost converges when using activations from approximately 1500 neurons, while activations from as few as 150 neurons yield relatively decent results across all test sets. Consequently, we provide detailed results for using 150 and 1500 neurons in Table 8. The table shows that, on average, safety neurons outperform other baselines, especially when fewer neurons are used. Additionally, the neurons from the last layer do not encode more information than neurons in various layers. These results suggest that the activations of safety neurons indeed encode more information about the safety of the model’s outputs, and this information is transferable across different datasets. The results of using this classifier as a safeguard for LLMs to reject unsafe responses are shown in Figure 13, from which we observe a consistent improvement on both SFT and DPO.
Figure 11: The accuracy on 4 different test datasets using different numbers of neuron activations from Llama2-7b. For safety neurons, we use those with the highest change scores.
(a) Mistral-7b
(b) Gemma-7b
(a) Llama2-7b
(b) Gemma-7b
Figure 13: Cost scores of Llama2-7b and Gemma-7b with safeguard trained on safety neuron activation collected on HH-Harmless and evaluated on various red-teaming benchmarks.
Model Beavertails RedTeam HarmBench JailbreakLLMs
Compare SFT with DPO on SFT Generation
Base $-7.16$ $-5.46$ $-4.73$ $-8.28$
SFT $\mathbf{-12.02}$ $\mathbf{-12.24}$ $\mathbf{-7.96}$ $\mathbf{-7.56}$
Compare SFT with DPO on DPO Generation
Base $-7.89$ $-6.96$ $-3.92$ $-8.57$
SFT $-11.80$ $-11.88$ $-7.87$ $-7.29$
Compare Base with DPO on Base Generation
Base $\mathbf{-10.41}$ $\mathbf{-9.51}$ $\mathbf{-7.38}$ $\mathbf{-9.11}$
SFT $-10.29$ $-10.90$ $-5.36$ $-5.13$
Compare Base with DPO on DPO Generation
Base $-9.15$ $-7.71$ $-3.00$ $-8.39$
SFT $-10.56$ $-11.11$ $-7.26$ $-6.44$
Table 9: The cost scores of Llama2-7b Base and Llama2-7b SFT patched with 20000 neurons’ activations from Llama2-7b DPO. The neurons are found via activation comparison from different models and generations. Bold denotes the best performance for Base and SFT respectively.
Model Beavertails RedTeam HarmBench JailbreakLLMs
Safety Neurons
Base $\mathbf{-7.16}$ $-5.46$ $\mathbf{-4.73}$ $\mathbf{-8.28}$
SFT $\mathbf{-12.02}$ $\mathbf{-12.24}$ $-7.96$ $\mathbf{-7.56}$
Prompt (all tokens)
Base $-4.19$ $\mathbf{-6.30}$ $-2.62$ $-1.19$
SFT $-9.23$ $-9.27$ $-2.95$ $-3.09$
Prompt (last token)
Base $-1.77$ $-0.83$ $4.51$ $-1.52$
SFT $-11.34$ $-11.61$ $\mathbf{-7.99}$ $-6.91$
Table 10: The cost scores of Llama2-7b Base and Llama2-7b SFT patched with 20000 neurons’ activations from Llama2-7b DPO. The neurons are found via activation comparison from different token positions. Bold denotes the best performance for Base and SFT respectively.

### v2

@@ E.4 More Alignment Tax Results
Spearman’s rank correlation coefficients between different preference neurons of Mistral-7b and Gemma-7b are shown in Figure 11. For Mistral-7b, we observe results similar to Llama2-7b. However, Gemma-7b shows anomalies when aligned on RewardBench-Safety, which we attribute to the small dataset size (less than 1k samples) compared to the larger number of neurons Gemma-7b. This discrepancy likely leads to insufficient training. However, this discrepancy does not affect our explanation of the alignment tax (Table 3).
Our experimental results indicate that when performing either safety or helpfulness alignment individually, safety neurons can be leveraged to improve both model safety and helpfulness. However, when both alignments are applied simultaneously, it remains unclear whether conflicts arise within these neurons. To investigate this, we conducted the following experiment on Llama2:
  • We used DPO to train two models based on the same SFT model: one trained solely on HH-helpful data (Helpful DPO) and another trained on both HH-harmless and HH-helpful data (HH DPO).
  • We patched $5\%$ of neuron activations from HH DPO into Helpful DPO, specifically using the neurons identified from Helpful DPO for model helpfulness. These neurons were selected following the same methodology used for identifying safety neurons in our study.
The results (Table 10) indicate that the neurons identified as helpful neurons are also crucial for improving model safety during HH training. This suggests that safety and helpfulness compete for control over these neurons during HH training, providing further mechanistic insights into the alignment tax phenomenon.
(a) Mistral-7b
(b) Gemma-7b
Table 10: Cost scores of Llama2 DPO models. Abbr. BT = Beavertails, RT = RedTeam, HB = HarmBench, JL = JailBreakLLMs, denotes patching helpfulness neurons’ activations from HH DPO.
BT RT HB JL
Helpful DPO 3.42 0.65 6.68 6.66
Helpful DPO -11.77 -11.09 -5.57 -8.28
HH DPO -11.81 -12.42 -10.41 -11.76
E.5 More Safeguard Results
@@
To validate the generalization ability of these neuron activations, we use activations from one dataset as the training set and merge the others as the test set, training a simple logistic regression classifier. Finally, we compute the average accuracy across all possible combinations as the evaluation metric. In addition to safety neurons, we employ neurons identified through other strategies as baselines, including (1) RN-Same Distribution, which refers to randomly sampled neurons (completely disjoint from safety neurons) with the same per-layer neuron count as the safety neurons; (2) RN-Last, which denotes neurons randomly sampled exclusively from the last layer, based on the hypothesis that neurons in the last layer directly influence the model’s output, making this a potentially strong baseline; (3) RN-All, which refers to neurons randomly sampled without constraints, aiming to assess whether the layer-wise distribution of safety neurons inherently encodes safety-related information. For all experiments requiring randomly sampled neurons, we repeat the process 5 times using different random seeds and report the averaged results.
@@
We train and test the classifier using activations from different numbers of neurons, as shown in Figure 10(b). The results indicate that the test accuracy almost converges when using activations from approximately $1500$ neurons, while activations from as few as $150$ neurons yield relatively decent results across all test sets. These results suggest that the activations of safety neurons indeed encode more information about the safety of the model’s outputs, and this information is transferable across different datasets. Additionally, random neurons with the same layer distribution as safety neurons are more effective than those sampled from other layers, which indicates the layer distribution of safety neurons may also encode safety information.
Appendix F Limitations
Appendix G Broader Impacts
This work contributes to the understanding of safety alignment in LLMs, potentially guiding the development of safer and more controllable models. By providing insights into the mechanisms behind safety neurons, our research may help improve alignment techniques and safeguard applications against generating harmful content. However, these findings could also be misused to weaken alignment constraints, enabling more covert adversarial attacks. Additionally, the reliance on specific datasets and evaluation methods may introduce biases that impact generalizability. To mitigate these risks, we adhere to open-source licensing policies and commit to sharing our code and data to foster transparency and responsible research.